OpenStudy (anonymous):

find the value of x or y so the line passing through the two points has the given slope (-3,y),(-9,-2); m = 1

OpenStudy (anonymous):

ready?

OpenStudy (anonymous):

since we know a line has the form y=mx+b, we can use the point-slope formula with the 2nd coordinate.

OpenStudy (anonymous):

y-y1 = m(x-x1) right?

OpenStudy (anonymous):

so now insert the coordinates (-9, -2) into that formula: y-(-2) = 1(x-(-9) or y +2 = x+9

OpenStudy (anonymous):

solve for y. can u do the rest?

OpenStudy (anonymous):

y = 4??

OpenStudy (anonymous):

y +2 = x+9 ... so move the 2 over by subtracting 2 from both sides to isolate y. then your answer will be in y=mx+b form

OpenStudy (anonymous):

y=x+7?

OpenStudy (anonymous):

yes!!